MONT 104N – Modeling the Environment Second “Debates” Day November 14, 2012

Presenters

Robertho Gay, Bobby Indeglia, Mike Keane, Marshall Lougee, Nick Lovullo, Joe Mc-Namara

Priorities for President Obama’s Second Term

Tuesday’s national election led to only slight rearrangements of the balance between the Republicans and the Democrats in Congress and President Obama narrowly won a second term by defeating Governor Romney. There will be no “honeymoon” period this time (as new Presidents sometimes get), though, because even before the new Congress takes office in January 2013, averting the impending “fiscal cliff” (the deep automatic cuts in most categories of Federal spending and across-the-board expiration of the Bush-era tax rate decreases) will require bipartisan action. Looking past that deadline, what should be the goals for the next four years?

There is an interesting collection of positions staked out in the New York Times Room For Debate blog posted on Wednesday, November 7:

              http://www.nytimes.com/roomfordebate/2012/11/07
            /a-to-do-list-for-president-obamas-next-four-years

(this should be all entered as one web address; it’s broken over two lines to fit on the page here) including several touching directly on issues we are considering in this course.

Assignment

For this “debate,” I think it will be a good plan for each of you to take the position of one (or maybe two) of the eight contributors to the New York Times blog, look into who your people are, what their backgrounds are, why they might hold the opinions they do, and then research some of the details behind what they are saying. Be prepared to discuss why your person (or people) think their issues are the most important ones that need to be addressed in President Obama’s second term. Look for connections with things we have discussed in class and highlight them in your presentations if appropriate.

Every can use the computer and projector in our classroom to display any graphical or numerical information they would like to bring in.

The natural way to do this will be to have each of you take one position and plan for each person to speak for about 4 or 5 minutes, with time for a follow-up discussion after everyone has spoken. If you want, I can assign who takes which position. But you can also sort that out for yourselves if you have strong feelings one way or another and you want to present your own opinions.

I would like to meet with the whole team before class on Wednesday, November 14 to see where you are and see if you have any questions about how to make your cases. I’m free basically during office hours 10am - 12 noon on Monday, November 12 and all morning on Tuesday, November 13.